a rectangular auditorium seats 2184 people.The number of seats in each row exceeds the number of rows by 17. find the number of seats in each row
Answer by fcabanski(1391)

Call the number of rows x


Then the number of seats in each row is x+17 (it exceeds the number of rows by 17).


The total number of seats, 2184 = the number of rows times the number of seats per row, which is x(x+17).


x(x+17)=2184 (subtract 2184 from both sides and multiply what's inside the () by x)


x&2 + 17x -2184 =0 (reverse foil )


(x+56)(x-39)=0 (solve each factor for x


x+56=0 (subtract 56 from both sides) x=-56. Throw it away because there can't be a negative number of rows.


x-39=0 (Add 39 to both sides.) x=39.
There are 39 rows, and 39+17 = 56 seats per row.